Solve the system of equations
x+y=6
x-y=-7
Set the first equation to: x=6-y, and solve the second equation:
(6-y)-y=-7
-y-y=-7-6
-2y=-13
y=13/2=6.5
Now solve the first equation with the newly found y value of 13/2 to obtain x:
x+(13/2)=6
x=6-(13/2)
x=-1/2=-0.5
You can check the answer by verifying if the two values satisfy either equation:
-0.5+6.5=6, which is correct and -0.5-6.5=7, which is correct as well.

Answer by student100(1)

Set the first equation to: x=6-y, and solve the second equation:
(6-y)-y=-7
-y-y=-7-6
-2y=-13
y=13/2=6.5
Now solve the first equation with the newly found y value of 13/2 to obtain x:
x+(13/2)=6
x=6-(13/2)
x=-1/2=-0.5
You can check the answer by verifying if the two values satisfy either equation:
-0.5+6.5=6, which is correct and -0.5-6.5=7, which is correct as well.